OpenStudy (anonymous):

Can the formula be written in the form of an exponential function or a power function? If not, explain why the function does not fit either form. 2. q(x) = 5^(x^2)

OpenStudy (anonymous):

sorry i glitched

OpenStudy (anonymous):

a formula can be used for a power function

OpenStudy (anonymous):

other then that idk

OpenStudy (anonymous):

i know there are formulas for each but I don't know how to get into the correct form

OpenStudy (mathmale):

You've typed in: q(x) = 5^(x^2) which can be written using Equation Editor (below) as \[q(x) = 5^{x ^{2}}\]

OpenStudy (mathmale):

First, Peter, from which course does this problem come? Are you familiar with power series?

OpenStudy (anonymous):

it is from the functions modeling change book 4th addition

OpenStudy (anonymous):

and yes I know what the power function looks like

OpenStudy (mathmale):

No, Peter, not power function, but power series. Familiar?

OpenStudy (anonymous):

i guess not then

OpenStudy (mathmale):

It just so happens that e^x can be approximated by a power series which consists of a long sum of powers of x. In other words, e^x is approx. equal to 1 +x/1 + x^2/2! + ... Have you seen this form before?

OpenStudy (anonymous):

No I have not

OpenStudy (mathmale):

My point is that your 5^(x^2) can be re-written as an exponential function which in turn can be approximated by a power series, which in turn consists of a long sum containing the first, the second, the third, (etc.) powers of x. But if you haven't studied power series yet, it's not practical to discuss this approach.

OpenStudy (mathmale):

Can you, Peter, rewrite the given function as an exponential function? Hint: \[5=e ^{\log_{e}5 }\]

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

\[\ln5 = \log _{e}5\] thats as far as i can get im not sure

OpenStudy (mathmale):

5^(x^2) can be re-written as \[(e ^{\ln 5})^{x ^{2}}=e ^{x ^{2}\ln 5}\]

OpenStudy (mathmale):

I'm unsure of what to tell you next, because my suggestion here depends upon our understanding of power series expansions. Unless I'm sorely mistaken, Peter, the last expression I've typed can be expanded in a power series.

OpenStudy (anonymous):

that will be fine thank you for the help!

OpenStudy (mathmale):

You're welcome. If you wait a while, this question of yours may be bumped to the top of the queue and someone else may offer to help you. Best wishes to you.

OpenStudy (mathmale):

Want me to bump you to the top of the queue now?

OpenStudy (anonymous):

sure

OpenStudy (mathmale):

OK!